From GlueXWiki
Jump to: navigation, search

July 5, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
    • Production status
    • Fourth package tests
    • Refurbishment preparations
  2. Engineering update
  3. Electronics update (Chris)
  4. LV cable assignment (Vlad)
  5. Test set-up in 126 (Beni)
  6. Other


Participants: Bill, Glen, Eugene, Chris, Nick, Simon, Vlad, Beni, and Lubomir


- Dave was absent but he updated the production tracking linked above: last two spare cathodes are almost ready. Spare wire frames: third wire frame (#6) is being tested already for a week, the current is very high (~1100nA) which doesn't surprise anybody since the caps were not replaced. Bill wanted to know if we have to order additional g10 frames. We have one spare frame now that will be used to replace wire frame #1. First we will try to fix #6, if it doesn't work we have to order g10 frames. The problem with #6 is that the caps are covered with Himiseal. Bill suggested, first, removing the caps (say on one HV sector) and then cleaning the Humiseal. Chirs: aceton cleans Humiseal very well (we can't use it inside the chamber!), but one can also use alcohol; have to soak the Humiseal first and after 5min start cleaning. We will try that.

- Started preparations of the first package (which is at Blue Crab) for refurbishment: HV cables from the PCBs were disconnected from the cables that are in turn connected to the two HV connectors. Later we may have problems to reconnect them since the cables were cut to the proper length, we have to route them exactly as they were before. Copper tubes (together with one part of the pre-amp brackets that is glued to the tube) and the pre-amps were removed from the package.

- Fourth package tests: Vlad tested cell #4 which is OK. On cell #3 (with wire frame #27) the current was very high (~800-1000nA) and doesn't go down already four days, but never tripped: most likely the HV caps have problems. Other than that all the channels on #3 are fine. Vlad applied today HV on cell #2 (wire frame #28), the current was initially 500nA but then started going up. It turned out Anatoly doesn't have a record that he had replaced the HV caps on #28. We will finished with the tests on cell #2 and cell #1 and then decide what exactly we want to disassemble.

Engineering update

- Bill gave us some detail about the cooling system: ~2gal, 32psi, tubes outside of the packages are (special) plastic 3/4" ID, brass compression fittings, manifolds are made out of Norel, tubes on the packages are copper 2.79mm ID. Water or Fluorinert: Fluorinert is $500/gal but volume is only 2 gal; the main concern is about leakages and special pump and fittings that are needed. Water leakage can cause damages, need to cover BCAL and try somehow to cover the manifolds. It turned out at the LHC experiments there's only one example when water is used for cooling a detector that is not accessible: the ECAL for CMS. In this case, however, many additional systems are included (see text above) to monitor the pressure, temperature and flow and in case of leakage to separate sections of the system and evacuate the water from there. All at the meeting, including Bill and Eugene, think Fluorinert is a better choice.

- Bill went through the drawings linked above, explaining the cooling system. There are two loops, inner and outer. The outer uses water. There's a heat-exchanger between the two loops. Beni and Eugene asked about its reliability, what is the pressure difference in the inner and outer loops. The cooling system cart is supposed to be placed on the floor downstream after the magnet, but Eugene suggested to keep this space empty for the future Cherenkov detector.

Electronics update

- Chris: no grounding clips yet. The sole source PR about the PCBs was returned to Chris: not enough reasons for the sole source. Chris will ask Fernando for help in formulating the sole source argumentation.

LV cable assignment

- Vlad: LV scheme is linked above, at the end of the file the cables are shown with the corresponding pre-amps and their labels for the cathodes: four pre-amps per cables. For the wires the cables are identified by the range of the channels on the PCB to which the LV cable is connected. We will put these lables on both sides of the LV cables.

Test set-up at 126

- Beni: the configuration of the DAQ to be used with two new Linux ROCs is close to complete. Old TIs will be used for the FDC DAQ system.


- Beni: next Wednesday (7/11) at 10am there will be a meeting about the gas system.